# Graphon-Theoretic Approach to Central Limit Theorems for $\epsilon$-Independence

### v1

@@ 1. Introduction
Interestingly, graphon limits have been used in the context of random matrix theory to encode the evolution of variance profiles and a corresponding limit theorem for the empirical spectral distribution of non-homogeneous Wigner matrices established accordingly [33]. It is also worth noting that other notions of independence, such as mixtures of free, monotone, and boolean independence, have been explored, with a corresponding central limit theorem established (see [15] and references therein).
@@
    For every $k\geq 1$, $i\in[n]^{k}$ such that $i_{j_{1}}=i_{j_{2}}$ and $j_{1}<j_{2}$, if there exists $j_{1}<j_{3}<j_{2}$ with $(i_{j_{1}},i_{j_{3}})\in E$, then for any centered random variables $a_{j}\in\mathcal{A}_{i_{j}}$, $j\in[k]$, we have
@@
converge to a Gaussian (resp. semi-circle) random variable with mean zero and variance equal to the variance of $a$. The combinatorial proof of those results involves the understanding of pair partitions. While all pairings play a role in the classical CLT, only noncrossing pairings contribute to the free CLT. For a pair partition $\pi\in P_{2}(2p)$, we define its intersection graph $f_{\pi}$ whose vertex set consists of the blocks of $\pi$, and there exists an edge between two blocks if they cross. Noncrossing pair partitions are those whose intersection graph consists of isolated vertices (see Section 2 for more details).
@@
The class of universal laws appearing at the limit in Theorem 1.1 is quite vast, and closed under classical and free convolution (see Proposition 2.4). The measures $\mu_{w}$ are universal objects, always arising as the limiting measures in CLTs for $\epsilon$-independent variables. This universality stems from the fact that every graphon $w$ can be realized as the limit of some sequence of (random) graphs (see [20, Corollary 2.6]), making $\mu_{w}$ the limiting measure associated with $\epsilon$-independent random variables defined by those (random) graphs. The breadth of this class reflects the diversity of graphon limits, encompassing an extensive and exotic variety of distributions. We refer to Section 7 for examples. One particular example which represents a relatively well-studied class of distributions interpolating between the Gaussian and semi-circle distributions is the class of $q$-Gaussians for $q\in[0,1]$ [7], whose $2p$-th moments are given by
@@
Theorem 1.1 recovers these distributions and the corresponding central limit theorem [24, Corollary 3], as they coincide with $\mu_{w}$ for the constant graphon $w(x,y)=q$ for all $x\neq y$. It is worth noting that $q$-Gaussians represent in some sense an extremal case of Theorem 1.1 as every limiting law $\mu_{w}$ appearing in Theorem 1.1 is stochastically dominated (in terms of moments) by some $q$-Gaussian distribution. This is formalized in the following proposition.
@@ 2.2. Fock representation of $\mu_{w}$
is the canonic operator corresponding to $\varphi$. According to [6, Theorem 2.2] the operators $P^{(n)}$ are positive, thus $\langle,\rangle_{T_{w}}$ is positive definite. If $\|T_{w}\|<1$ then, by [6, Theorem 2.3], we have that all the $P^{(n)}$’s are strictly positive and we can take $\mathcal{F}_{T_{w}}$ as the completion of $\mathcal{F}$ with respect to $\langle,\rangle_{T_{w}}$. When $\|T_{w}\|=1$, the kernel might not be empty and the completion is made after taking the corresponding quotient.
@@
From [6, Theorem 3.1], we know that
@@ 2.3. First properties
It is a consequence (following [18, Proof of Theorem 4.1]) of the fact that we have multiplicativity of the term
@@
Let $n_{\lambda}=\lceil\lambda n\rceil$. By [1, Theorem 4.8], we can find two sequences of graphs $g_{n},g^{\prime}_{n}$ over $n_{\lambda}$ and $n-n_{\lambda}$ vertices, respectively, such that $g_{n}\to w$ and $g^{\prime}_{n}\to w^{\prime}$. We label their vertices as
@@
Proposition 2.4 can also be proved by a direct computation using the definitions of $w^{*}$ and $\overline{w}^{*}$ in (8) and (9). For the classical convolution, this computation is reminiscent of the proof of [17, Proposition 3.1].
@@
The proof follows the lines of that in [23] with a few adaptations, which we outline below while omitting the lengthy details. The main ingredient is the explicit construction of a dual system for $A(1)$ on the Fock space. As in [23], the finiteness of the free Fisher information is a consequence of the existence of this dual system, thanks to [27, Theorem 1]. Moreover, by [22, Proposition 8.18], the finiteness of the free Fisher information implies the absolute continuity with respect to the Lebesgue measure, and the fact that the density is in $L^{3}(\mathbb{R})$.
In order to define the dual operator of $A(1)$, we use below the notation of [23]. In particular, $B(n+1)$ denotes a collection of incomplete pairings of $\{n>n-1>\cdots>1>0\}$ introduced in [23, Section 4], where the rules of counting crossings differ from the usual definition of crossings in $P_{1,2}(n+1)$. The new intersection graph $\tilde{f}_{\pi}$ of a partition $\pi\in B(n+1)$ is the graph over the blocks of $\pi$ such that two blocks are connected if they cross according to these new rules. For $\pi\in B(n+1)$, we denote by $p(\pi)$ the set of pairings in $\pi$, by $s(\pi)$ the singletons in $\pi$ and by $\pi(0)$ the integer which is coupled to 00 in $\pi$.
Following [23, Proposition 4.5], we define an unbounded operator $D$ with the domain $\mathcal{F}=\bigoplus_{n=0}^{\infty}H^{\otimes n}$ by linear extension of $D\Omega=0$ and $Dh=\sum_{\pi\in B(n+1)}h_{\pi}$ whenever $h\in L^{2}([0,1]^{n})$, where for any $\pi\in B(n+1)$ with $\ell$ singletons and $m$ pairings, $h_{\pi}$ is the element of $L^{2}([0,1])^{\ell}$ given by
@@
(note that all the variables $x_{i}$ and $x_{j}$ are indexed by the blocks of $\tilde{f}_{\pi}$, and the blocks $\bar{n},\ldots,\bar{1}$ denote respectively the blocks of $n,\ldots,1$ in $\tilde{f}_{\pi}$). Using the explicit formula (6) of $c(h)^{*}$ and following the proof of [23, Proposition 4.5], we get that the operator $D$ satisfies
@@
where $P$ is the orthogonal projection onto $\mathbb{C}\Omega$. In order to conclude, it remains to prove that $\Omega$ is in the domain of $D^{*}$, and again, the proof follows the one of [23, Theorem 4.6]: it consists in showing that the linear functional $\langle D\ \cdot\ ,\Omega\rangle_{T_{w}}$ is bounded by computing $\langle Dh,\Omega\rangle_{T_{w}}$ for $h$ being any finite linear combination of pure vectors, and bounding it by a constant times $\|h\|_{T_{w}}$. One crucial ingredient in the proof is the fact that the free right annihilation operators are bounded on $\mathcal{F}=\bigoplus_{n=0}^{\infty}H^{\otimes n}$. This fact, which is the content of [23, Lemma 2.2], remains valid in our context since [4, Theorem 6], used to prove [23, Lemma 2.2], is stated in a level of generality which applies to our setting. ∎
@@
The following result provides the moment-cumulant formula for $\epsilon$-independence [29, Theorem 5.2].
@@
It was proved in [2] that $\rho(f,(g_{n},\alpha_{n}))$ converges if and only if $w_{g_{n},\alpha_{n}}$ is a Cauchy sequence in the metric $\delta_{\Box}$. In this case, there exists a limit $w$ of the graphons $w_{g_{n},\alpha_{n}}$ [2, Corollary 3.9] and we have
@@ 5. Multivariate case
the graphon $w_{G_{g_{n}}}$ is well-defined as an $\mathcal{M}_{L}$-decorated graphon. It was proved in [16, Theorem 3.7] that, given a sequence of $\mathcal{M}_{L}$-decorated graphons $(w_{n})_{n\geq 1}$, if $\rho(F,w_{n})$ converges as $n\to\infty$ for every $\mathcal{M}_{L}$-decorated graph $F$, then there exists an $\mathcal{M}_{L}$-decorated graphon $w$ such that $\rho(F,w_{n})\to\rho(F,w)$, for every $\mathcal{M}_{L}$-decorated graph $F$. We denote this convergence by $w_{n}\to w$. Let us note that it is possible to define the cut distance $\delta_{\Box}$ on the space of $\mathcal{M}_{L}$-decorated graphons in such a way that, for any simple graph $F$, the map $w\mapsto\rho(F,w)$ is continuous for $\delta_{\Box}$. More precisely, set
@@
This distance $\delta_{\Box}$ gives a criterion for the convergence $w_{n}\to w$. The proof of such continuity is carried out in [16, Section 2].
@@
By [16, Lemma 2.3] and a similar argument done in Lemma 4.1, we can bound
@@ 6. Graphon with value in $[-1,1]$
Note that in [28, Theorem 2], Speicher considered a sequence of symmetric $n\times n$ matrices with entries in $\{-1,1\}$ with independent and identically distributed variables (up to the symmetry). In this case, the law of large numbers implies that $w_{Q_{n}}$ converges almost surely in the graphon sense to a constant graphon, which allows us to recover [28, Theorem 2] by using Theorem 6.1.

### v2

@@ 1. Introduction
Interestingly, graphon limits have been used in the context of random matrix theory to encode the evolution of variance profiles and a corresponding limit theorem for the empirical spectral distribution of non-homogeneous Wigner matrices established accordingly [34]. It is also worth noting that other notions of independence, such as mixtures of free, monotone, and boolean independence, have been explored, with a corresponding central limit theorem established (see [1, 16] and references therein).
@@
    For every $k\geq 1$, $i\in[n]^{k}$ such that for all $j_{1}<j_{2}$ with $i_{j_{1}}=i_{j_{2}}$, there exists $j_{1}<j_{3}<j_{2}$ with $(i_{j_{1}},i_{j_{3}})\notin E$, then for any centered random variables $a_{j}\in\mathcal{A}_{i_{j}}$, $j\in[k]$, we have
@@
converge in distribution to a Gaussian (resp. semi-circle) random variable with mean zero and variance equal to the variance of $a$. The combinatorial proof of those results involves the understanding of pair partitions. While all pairings play a role in the classical CLT, only noncrossing pairings contribute to the free CLT. For a pair partition $\pi\in P_{2}(2p)$, we define its intersection graph $f_{\pi}$ whose vertex set consists of the blocks of $\pi$, and there exists an edge between two blocks if they cross. Noncrossing pair partitions are those whose intersection graph consists of isolated vertices (see Section 2 for more details).
@@
The class of universal laws appearing at the limit in Theorem 1.1 is quite vast, and closed under classical and free convolution (see Proposition 2.4). The measures $\mu_{w}$ are universal objects, always arising as the limiting measures in CLTs for $\epsilon$-independent variables. This universality stems from the fact that every graphon $w$ can be realized as the limit of some sequence of (random) graphs (see [21, Corollary 2.6]), making $\mu_{w}$ the limiting measure associated with $\epsilon$-independent random variables defined by those (random) graphs. The breadth of this class reflects the diversity of graphon limits, encompassing an extensive and exotic variety of distributions. We refer to Section 7 for examples. One particular example which represents a relatively well-studied class of distributions interpolating between the Gaussian and semi-circle distributions is the class of $q$-Gaussians for $q\in[0,1]$ [8], whose $2p$-th moments are given by
@@
Theorem 1.1 recovers these distributions and the corresponding central limit theorem [25, Corollary 3], as they coincide with $\mu_{w}$ for the constant graphon $w(x,y)=q$ for all $x\neq y$. It is worth noting that $q$-Gaussians represent in some sense an extremal case of Theorem 1.1 as every limiting law $\mu_{w}$ appearing in Theorem 1.1 is stochastically dominated (in terms of moments) by some $q$-Gaussian distribution. This is formalized in the following proposition.
@@ 2.2. Fock representation of $\mu_{w}$
is the canonic operator corresponding to $\varphi$. According to [7, Theorem 2.2] the operators $P^{(n)}$ are positive, thus $\langle,\rangle_{T_{w}}$ is positive definite. If $\|T_{w}\|<1$ then, by [7, Theorem 2.3], we have that all the $P^{(n)}$’s are strictly positive and we can take $\mathcal{F}_{T_{w}}$ as the completion of $\mathcal{F}$ with respect to $\langle,\rangle_{T_{w}}$. When $\|T_{w}\|=1$, the kernel might not be empty and the completion is made after taking the corresponding quotient.
@@
From [7, Theorem 3.1], we know that
@@ 2.3. First properties
It is a consequence (following [19, Proof of Theorem 4.1]) of the fact that we have multiplicativity of the term
@@
Let $n_{\lambda}=\lceil\lambda n\rceil$. By [2, Theorem 4.8], we can find two sequences of graphs $g_{n},g^{\prime}_{n}$ over $n_{\lambda}$ and $n-n_{\lambda}$ vertices, respectively, such that $g_{n}\to w$ and $g^{\prime}_{n}\to w^{\prime}$. We label their vertices as
@@
Proposition 2.4 can also be proved by a direct computation using the definitions of $w^{*}$ and $\overline{w}^{*}$ in (8) and (9). For the classical convolution, this computation is reminiscent of the proof of [18, Proposition 3.1].
@@
The proof follows the lines of that in [24] with a few adaptations, which we outline below while omitting the lengthy details. The main ingredient is the explicit construction of a dual system for $A(1)$ on the Fock space. As in [24], the finiteness of the free Fisher information is a consequence of the existence of this dual system, thanks to [28, Theorem 1]. Moreover, by [23, Proposition 8.18], the finiteness of the free Fisher information implies the absolute continuity with respect to the Lebesgue measure, and the fact that the density is in $L^{3}(\mathbb{R})$.
In order to define the dual operator of $A(1)$, we use below the notation of [24]. In particular, $B(n+1)$ denotes a collection of incomplete pairings of $\{n>n-1>\cdots>1>0\}$ introduced in [24, Section 4], where the rules of counting crossings differ from the usual definition of crossings in $P_{1,2}(n+1)$. The new intersection graph $\tilde{f}_{\pi}$ of a partition $\pi\in B(n+1)$ is the graph over the blocks of $\pi$ such that two blocks are connected if they cross according to these new rules. For $\pi\in B(n+1)$, we denote by $p(\pi)$ the set of pairings in $\pi$, by $s(\pi)$ the singletons in $\pi$ and by $\pi(0)$ the integer which is coupled to 00 in $\pi$.
Following [24, Proposition 4.5], we define an unbounded operator $D$ with the domain $\mathcal{F}=\bigoplus_{n=0}^{\infty}H^{\otimes n}$ by linear extension of $D\Omega=0$ and $Dh=\sum_{\pi\in B(n+1)}h_{\pi}$ whenever $h\in L^{2}([0,1]^{n})$, where for any $\pi\in B(n+1)$ with $\ell$ singletons and $m$ pairings, $h_{\pi}$ is the element of $L^{2}([0,1])^{\ell}$ given by
@@
(note that all the variables $x_{i}$ and $x_{j}$ are indexed by the blocks of $\tilde{f}_{\pi}$, and the blocks $\bar{n},\ldots,\bar{1}$ denote respectively the blocks of $n,\ldots,1$ in $\tilde{f}_{\pi}$). Using the explicit formula (6) of $c(h)^{*}$ and following the proof of [24, Proposition 4.5], we get that the operator $D$ satisfies
@@
where $P$ is the orthogonal projection onto $\mathbb{C}\Omega$. In order to conclude, it remains to prove that $\Omega$ is in the domain of $D^{*}$, and again, the proof follows the one of [24, Theorem 4.6]: it consists in showing that the linear functional $\langle D\ \cdot\ ,\Omega\rangle_{T_{w}}$ is bounded by computing $\langle Dh,\Omega\rangle_{T_{w}}$ for $h$ being any finite linear combination of pure vectors, and bounding it by a constant times $\|h\|_{T_{w}}$. One crucial ingredient in the proof is the fact that the free right annihilation operators are bounded on $\mathcal{F}=\bigoplus_{n=0}^{\infty}H^{\otimes n}$. This fact, which is the content of [24, Lemma 2.2], remains valid in our context since [5, Theorem 6], used to prove [24, Lemma 2.2], is stated in a level of generality which applies to our setting. ∎
@@
The following result provides the moment-cumulant formula for $\epsilon$-independence [30, Theorem 5.2].
@@
It was proved in [3] that $\rho(f,(g_{n},\alpha_{n}))$ converges if and only if $w_{g_{n},\alpha_{n}}$ is a Cauchy sequence in the metric $\delta_{\Box}$. In this case, there exists a limit $w$ of the graphons $w_{g_{n},\alpha_{n}}$ [3, Corollary 3.9] and we have
@@ 5. Multivariate case
the graphon $w_{G_{g_{n}}}$ is well-defined as an $\mathcal{M}_{L}$-decorated graphon. It was proved in [17, Theorem 3.7] that, given a sequence of $\mathcal{M}_{L}$-decorated graphons $(w_{n})_{n\geq 1}$, if $\rho(F,w_{n})$ converges as $n\to\infty$ for every $\mathcal{M}_{L}$-decorated graph $F$, then there exists an $\mathcal{M}_{L}$-decorated graphon $w$ such that $\rho(F,w_{n})\to\rho(F,w)$, for every $\mathcal{M}_{L}$-decorated graph $F$. We denote this convergence by $w_{n}\to w$. Let us note that it is possible to define the cut distance $\delta_{\Box}$ on the space of $\mathcal{M}_{L}$-decorated graphons in such a way that, for any simple graph $F$, the map $w\mapsto\rho(F,w)$ is continuous for $\delta_{\Box}$. More precisely, set
@@
This distance $\delta_{\Box}$ gives a criterion for the convergence $w_{n}\to w$. The proof of such continuity is carried out in [17, Section 2].
@@
By [17, Lemma 2.3] and a similar argument done in Lemma 4.1, we can bound
@@ 6. Graphon with value in $[-1,1]$
Note that in [29, Theorem 2], Speicher considered a sequence of symmetric $n\times n$ matrices with entries in $\{-1,1\}$ with independent and identically distributed variables (up to the symmetry). In this case, the law of large numbers implies that $w_{Q_{n}}$ converges almost surely in the graphon sense to a constant graphon, which allows us to recover [29, Theorem 2] by using Theorem 6.1.
@@ 7. Examples
To conclude, we give some nontrivial measures associated with negative graphons.
Example 7.5.
The following laws
[…]
Proof.
The first case follows from the fact that the Rademacher law is the $q$-Gaussian law $\mu_{q}$ for $q=-1$. Note that the arcsine law can be written as $\frac{1}{\sqrt{2}}(\mu_{-1}\boxplus\mu_{-1})$, and more generally, the Kesten-McKay law can be written as $\frac{1}{\sqrt{d}}(\mu_{-1})^{\boxplus d}$; see [27, Example 12.8, Exercise 12.21]. We conclude by Proposition 2.4 and Remark 2.5. ∎